# Are Self-Attentions Effective for Time Series Forecasting?

### v1

@@ 4 Experiments
In this section, we provide extensive experiments to provide the benefits of our proposed framework, CATS, including forecasting performance and computational efficiency. To this end, we use 7 different real-world datasets and 9 baseline models. For datasets, we use Electricity, ETT (ETTh1, ETTh2, ETTm1, and ETTm2), Weather, and Traffic. These datasets are provided in Wu et al. (2021) for time series forecasting benchmark, detailed in Appendix.
For baselines, we utilize a wide range of various baselines, including the state-of-the-art long-term time series forecasting model TimeMixer (Wang et al., 2024), PatchTST (Nie et al., 2023), Timesnet (Wu et al., 2022), Crossformer (Zhang and Yan, 2022), MICN (Wang et al., 2022), FiLM (Zhou et al., 2022a), DLinear (Zeng et al., 2023), Autoformer (Wu et al., 2021), and Informer (Zhou et al., 2021). We used 4 NVIDIA RTX 4090 24GB GPUs with 2 Intel(R) Xeon(R) Gold 5218R CPUs @ 2.10GHz for all experiments.
@@ 4.1 Long-Term Time Series Forecasting Results
To ease comparison, we follow the settings of the most recent work Wang et al. (2024) for long-term forecasting, which uses various horizon lengths with fixed 96 input sequence lengths. The detailed experimental settings can be found in Appendix. Table 2 summarizes the forecasting performance for all datasets and baselines. Our proposed model, CATS, demonstrates superior performance across multiple datasets in multivariate long-term forecasting tasks. Specifically, CATS consistently achieves the lowest Mean Squared Error (MSE) and Mean Absolute Error (MAE) on the Weather and Electricity datasets, with MSE values of 0.161 and 0.149 for the 96-step horizon, respectively, outperforming all other models. Additionally, for the Traffic dataset, CATS shows competitive performance, resulting in the second-best results with an MSE of 0.421 and MAE of 0.270 for the 96-step horizon. This indicates that CATS effectively captures the underlying patterns in diverse types of time series data. We observe consistent state-of-the-art performance of CATS in terms of error metrics, which highlights the high performance of our model in handling complex temporal dependencies. We also provide additional experiments with longer 512 input sequence lengths in Appendix.
@@ 4.2 Efficient and Robust Forecasting for Long Input Sequences
Zeng et al. (2023) observed that many models experience a decline in performance when using long input sequences for time series forecasting. To address this, some approaches have been developed to capture long-term dependencies. For instance, TimeMixer Wang et al. (2024) employs linear models with mixed scale, and PatchTST Nie et al. (2023) utilizes an encoder network to encode long-term information. However, these models still have computational issues, particularly in terms of escalating memory and parameter requirements. Thus, in this subsection, we provide a comparison between previous models and ours in terms of efficient and robust forecasting for long input sequences.
@@
Table 3: Comparison of models with the number of parameters, GPU memory consumption, and MSE across different input sequence lengths on ETTm1. Full results with more diverse input lengths are presented in Appendix.
@@
Furthermore, we conduct a deeper comparison between transformer-based models. Especially, TimeMixer Wang et al. (2024) argues that their model outperforms PatchTST Nie et al. (2023) in the setting of long input sequences. Regarding this setting, we also conduct an experiment on $L=512$. We summarize the results in Fig. 4. Among these transformer-based models, our model achieves the lowest MSE for most forecasting horizons. Moreover, our model requires even a lower number of parameters, GPU memory, and running time. Especially, for parameter efficiency, CATS shows significant differences even on a log scale due to its efficient parameter-sharing. Fig. 4c highlights GPU memory usage across different forecasting horizons. While PatchTST and TimeMixer consume significantly more memory, CATS maintains a low and stable memory consumption, demonstrating superior memory efficiency. In Fig. 4d CATS also consistently achieves lower running times compared to PatchTST and TimeMixer.
@@ 4.3 Replacement of Cross-attention with Self-attention
4.4 Understanding Periodic Patterns with Cross-Attention
As noted in Section 3.2, in our proposed model, all operations including embedding and multi-head attention are performed independently for each learnable query. In other words, the forecast for a specific horizon does not depend on other horizons. This approach allows us to gain a clear understanding of how each prediction is calculated. Therefore, in this subsection, we visualize how the proposed model understands the periodic properties.
@@ 4.4 Understanding Periodic Patterns with Cross-Attention
Fig. 6 illustrates (a) forecasting results, (b) a cross-attention score map (5$\times$4) on the ETTm1 dataset, and (c, d) the two pairs with the highest attention scores. We predict 96 steps with input sequence length 96 on ETTm1. The input patches consist of four patches of 24 lengths and one padding patch. As shown in Fig. 6c and 6d, the patches with high attention scores exhibit similar temporal patterns, demonstrating the ability of CATS to detect sequential and periodic patterns.
@@ 5 Conclusion
Based on our study, we exploit the advantages of Transformer models in time series forecasting by removing self-attentions and developing a new cross-attention-based architecture. We believe that our model establishes a strong baseline for such forecasting tasks and offers further insights into the complexities of long-term forecasting problems. Our findings provide a reevaluation of self-attentions in this domain, and we hope that future research can critically assess the efficacy and efficiency across various time series analysis tasks. As a limitation, our proposed methods assume channel independence between variables based on the recent work Nie et al. (2023). As the time series data in the real-world are highly correlated, we hope future research can address cross-variate dependency with reduced computation complexity based on the proposed architecture.
@@ A.1 Datasets
We evaluated the performance using seven datasets commonly used in long-term time series forecasting, including Weather, Traffic, Electricity, and ETT (ETTh1, ETTh2, ETTm1, and ETTm2). These datasets capture a range of periodic characteristics and scenarios that are difficult to predict in the real world, making them highly suitable for tasks, such as long-term time series forecasting, generation, and imputation. Details of these datasets are described in Table 5. These datasets are provided in Wu et al. [2021].
Table 5: Details of 7 real-world datasets.
Dimension Frequency Timesteps Information
Weather 21 10-min 52,696 Weather
Electricity 321 Hourly 17,544 Electricity
Traffic 862 Hourly 26,304 Transportation
ETTh1 7 Hourly 17,420 Temperature
ETTh2 7 Hourly 17,420 Temperature
ETTm1 7 15-min 69,680 Temperature
ETTm2 7 15-min 69,680 Temperature
@@ A.2 Hyperparameter Settings
In every experiment in our paper, following Nie et al. [2023], we fixed the random seed of 2021 to enhance the reproducibility of our experiments. Additionally, following numerous studies in the field of time series forecasting Nie et al. [2023], we fixed the input sequence length $L=96$. For the forecasting horizon $T$, we also used the widely accepted values, i.e., [96,192,336,720][96,192,336,720][ 96 , 192 , 336 , 720 ]. For our model, in all configurations, we adopt the GeGLU activation function [Shazeer, 2020] between the two linear layers in the feed-forward network for our model. Additionally, we use learnable positional embedding parameters for the input data and omit positional embeddings for learnable queries to avoid redundant parameter learning. For Table 2, our model uses three cross-attention layers with embedding size $D=256$, number of attention heads $H=32$. Specifically, to avoid overfitting on small datasets Nie et al. [2023], we use patch length 48 on the ETTh1 and ETTh2 datasets. Table 6 shows further details of hyperparameter settings.
Table 6: Experimental settings mainly used for the main paper.
@@ B.1 Performance with Longer Input Sequences
In Table 3, we compared models with the number of parameters, GPU memory consumption, and MSE across different input lengths on ETTm1 with varying input sequence lengths. In this section, we provide comprehensive results on longer input sequence lengths $L=512$. The detailed parameters can be found in Table 7 and the corresponding experimental results are summarized in Table 8. As with unified hyperparameter settings, we follow the settings of the most recent work [Wang et al., 2024] to ease comparison. Overall, the experimental results clearly illustrate the superiority of CATS over recent forecasting models across multiple datasets and prediction horizons. CATS consistently shows the lowest Mean Squared Error (MSE) and Mean Absolute Error (MAE) across a variety of datasets and forecast horizons. For instance, on Electricity, at the 96 forecast horizon, CATS achieves the best MSE score of 0.144 and the best MAE score of 0.189, underscoring its accuracy in predicting electrical demand.
@@
We provide additional experimental results to support the findings discussed in Section 4.2. Tables 9 and 10 summarize detailed comparisons of the number of parameters, GPU memory consumption, and MSE across different input lengths for the ETTm1 and Weather datasets, respectively. The linear models, TimeMixer and DLinear, exhibit smaller parameters for shorter input lengths. Despite CATS having slightly more parameters than TimeMixer for smaller inputs, it outperforms in terms of memory usage and MSE. This suggests that CATS is more efficient and effective in handling shorter inputs. For PatchTST, the number of parameters does not increase within the actual Transformer backbone as the input length increases. However, due to the need to flatten and project all inputs at the end, the parameters scale linearly with the input length. This highlights a limitation of the Encoder’s architecture. On the other hand, TimeMixer’s parameters grow almost quadratically as the input length doubles. Similarly, DLinear’s parameters increase linearly with the input length. Our proposed model, CATS demonstrates significant efficiency through parameter sharing, where the parameters hardly increase with longer inputs. Notably, from an input length of 336, CATS has fewer parameters than DLinear, showcasing the deep learning model’s advantage in detecting inherent patterns in the data.
@@ B.3 Ablation Study on Query-adaptive Masking
Figure 7: Comparison of performance with query-adaptive masking with two different probabilities, dropout, and using both query-adaptive masking and dropout. The results of $p=0.1$ to 0.70.70.7 indicate a probability masking that is linearly increased proportionally to the horizon predicted by the query.
@@
To push further, we reproduce the experiment of Fig. 6 with other datasets used in forecasting tasks. We illustrate the results of the Weather, Traffic, Electricity, ETTm2, ETTh1, and ETTh2 in Figures 9, 10, 11, 12, 13, and 14, respectively. For each figure, (a) represents the forecasting results, (b) shows the cross-attention score map, and (c) and (d) illustrate the two pairs with the highest attention scores. For all figures, our attention-based explanation successfully discovers similar periodic patterns. Therefore, we believe that our model has the potential to provide a clearer understanding of the mechanisms underlying forecasting predictions. We hope that future research will continue to explore and expand upon this foundation.

### v3

@@ 5 Experiments
In this section, we provide extensive experiments to provide the benefits of our proposed framework, CATS, including forecasting performance and computational efficiency. To this end, we use 7 different real-world datasets and 9 baseline models. For datasets, we use Electricity, ETT (ETTh1, ETTh2, ETTm1, and ETTm2), Weather, Traffic, and M4. These datasets are provided in [23] and [24] for time series forecasting benchmark, detailed in Appendix.
For baselines, we utilize a wide range of various baselines, including the state-of-the-art long-term time series forecasting model TimeMixer [21], PatchTST [14], Timesnet [24], Crossformer [28], MICN [20], FiLM [30], DLinear [26], Autoformer [23], and Informer [29]. For both long-term and short-term time series forecasting results, we report performance of our model alongside the results of other models as presented in TimeMixer [21], ensuring a consistent comparison across all baselines. We used 4 NVIDIA RTX 4090 24GB GPUs with 2 Intel(R) Xeon(R) Gold 5218R CPUs @ 2.10GHz for all experiments.
@@ 5.1 Long-Term Time Series Forecasting Results
To ease comparison, we follow the settings of [21] for long-term forecasting, using various forecast horizons with a fixed 96 input sequence length. Detailed settings are provided in the Appendix. Table 9 summarizes the forecasting performance across all datasets and baselines. Our proposed model, CATS, demonstrates superior performance in multivariate long-term forecasting tasks across multiple datasets. CATS consistently achieves the lowest Mean Squared Error (MSE) and Mean Absolute Error (MAE) on the Traffic dataset for all forecast horizons, outperforming all other models. For the Weather, Electricity, and ETT datasets, CATS shows competitive performance, achieving the best results on most forecast horizons. This indicates that CATS effectively captures underlying patterns in diverse time series data, highlighting its capability to handle complex temporal dependencies. Additional experiments with longer input sequence lengths of 512 are provided in the Appendix.
Table 2: Multivariate long-term forecasting results with recent forecasting models and ours for unified hyperparameter settings. The best results are in bold and the second best are underlined. Full results are provided in Appendix.
@@ 5.2 Efficient and Robust Forecasting for Long Input Sequences
Zeng et al. [26] observed that many models experience a decline in performance when using long input sequences for time series forecasting. To address this, some approaches have been developed to capture long-term dependencies. For instance, TimeMixer [21] employs linear models with mixed scale, and PatchTST [14] utilizes an encoder network to encode long-term information. However, these models still have computational issues, particularly in terms of escalating memory and parameter requirements. Thus, in this subsection, we provide a comparison between previous models and ours in terms of efficient and robust forecasting for long input sequences.
@@
Table 3: Comparison of models with the number of parameters, GPU memory consumption, and MSE across different input sequence lengths on ETTm1. Full results with more diverse input lengths are provided in Appendix.
@@
Furthermore, we conduct a deeper comparison between transformer-based models. Especially, TimeMixer [21] argues that their model outperforms PatchTST [14] in the setting of long input sequences. Regarding this setting, we also conduct an experiment on $L=512$. We summarize the results in Fig. 5. Among these transformer-based models, our model achieves the lowest MSE for most forecasting horizons. Moreover, our model requires even a lower number of parameters, GPU memory, and running time. Especially, for parameter efficiency, CATS shows significant differences even on a log scale due to its efficient parameter-sharing. Fig. 5c highlights GPU memory usage across different forecasting horizons. While PatchTST and TimeMixer consume significantly more memory, CATS maintains a low and stable memory consumption, demonstrating superior memory efficiency. In Fig. 5d CATS also consistently achieves lower running times compared to PatchTST and TimeMixer.
@@
5.3 Short-Term Time Series Forecasting Results
Table 4 summarizes the averaged results for the M4 dataset, comparing the performance of various models. Our CATS model consistently achieved the best results across all metrics. Notably, CATS reduced MASE by 19.94% compared to PatchTST, a self-attention-only model that suffers from temporal information loss as mentioned in Section 3. CATS, with its cross-attention-only structure, effectively mitigated this issue and captured temporal dependencies more efficiently than previous models. Although TimeMixer, a state-of-the-art linear model, performed well, CATS surpassed it across all metrics. This demonstrates that CATS excels at capturing short-term temporal dependencies, providing superior performance in short-term forecasting tasks.
Table 4: Averaged univariate short-term forecasting results in the M4 dataset. The best results are in bold and the second best are underlined. Full results are presented in Appendix.
Models CATS TimeMixer Timesnet PatchTST MICN FiLM DLinear Autoformer Informer
Average SMAPE 11.701 11.723 11.829 13.152 19.638 14.863 13.639 12.909 14.086
MASE 1.557 1.559 1.585 1.945 5.947 2.207 2.095 1.771 2.718
OWA 0.838 0.840 0.851 0.998 2.279 1.125 1.051 0.939 1.230
@@ 5.4 Replacement of Cross-attention with Self-attention
5.5 Explaining Periodic Patterns through Cross-Attention
As noted in Section 4.2, in our proposed model, all operations including embedding and multi-head attention are performed independently for each learnable query. In other words, the forecast for a specific horizon does not depend on other horizons. This approach helps us better understand how each prediction is derived. Therefore, in this subsection, we visualize how the proposed model understands the periodic properties.
@@ 5.5 Explaining Periodic Patterns through Cross-Attention
Fig. 7 illustrates (a) forecasting results, (b) a cross-attention score map (5$\times$4) on the ETTm1 dataset, and (c, d) the two pairs with the highest attention scores. We predict 96 steps with input sequence length 96 on ETTm1. The input patches consist of four patches of 24 lengths and one padding patch. As shown in Fig. 7c and 7d, the patches with high attention scores exhibit similar temporal patterns, demonstrating the ability of CATS to detect sequential and periodic patterns.
@@ 6 Conclusion
Based on our study, we exploit the advantages of Transformer models in time series forecasting by removing self-attentions and developing a new cross-attention-based architecture. We believe that our model establishes a strong baseline for such forecasting tasks and offers further insights into the complexities of long-term forecasting problems. Our findings provide a reevaluation of self-attentions in this domain, and we hope that future research can critically assess the efficacy and efficiency across various time series analysis tasks. As a limitation, our proposed methods assume channel independence between variables based on the recent work [14]. As the time series data in the real-world are highly correlated, we hope future research can address cross-variate dependency with reduced computation complexity based on the proposed architecture.
7 Acknowledgements
This work was partly supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2022-II220984, Development of Artificial Intelligence Technology for Personalized Plug-and-Play Explanation and Verification of Explanation) and the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. RS-2024-00338859). This work was also supported by the MSIT(Ministry of Science and ICT), Korea, under the ITRC(Information Technology Research Center) support program (IITP-2024-RS-2024-00438056) supervised by the IITP.
@@ A.1 Datasets
We evaluated the performance using seven datasets commonly used in long-term time series forecasting, including Weather, Traffic, Electricity, ETT (ETTh1, ETTh2, ETTm1, and ETTm2), and M4. These datasets capture a range of periodic characteristics and scenarios that are difficult to predict in the real world, making them highly suitable for tasks, such as long-term time series forecasting, generation, and imputation. Details of these datasets are described in Table 6. The M4 dataset are provided Wu et al. [24], while the remaining datasets are provided in Wu et al. [23].
Table 6: Details of 13 real-world datasets.
Dimension Frequency Timesteps Information Forecasting Horizon
Weather 21 10-min 52,696 Weather (96, 192, 336, 720)
Electricity 321 Hourly 17,544 Electricity (96, 192, 336, 720)
Traffic 862 Hourly 26,304 Transportation (96, 192, 336, 720)
ETTh1 7 Hourly 17,420 Temperature (96, 192, 336, 720)
ETTh2 7 Hourly 17,420 Temperature (96, 192, 336, 720)
ETTm1 7 15-min 69,680 Temperature (96, 192, 336, 720)
ETTm2 7 15-min 69,680 Temperature (96, 192, 336, 720)
M4-Quartely 1 Quartely 48000 Finance 8
M4-Monthly 1 Monthly 96000 Industry 18
M4-Yearly 1 Yearly 46000 Demographic 6
M4-Weekly 1 Weekly 718 Macro 13
M4-Daily 1 Daily 8454 Micro 14
M4-Hourly 1 Hourly 828 Other 48
@@ A.2 Hyperparameter Settings
In every experiment in our paper, following [14], we fixed the random seed of 2021 to enhance the reproducibility of our experiments. Additionally, following numerous studies in the field of time series forecasting [14], we fixed the input sequence length $L=96$. For the forecasting horizon $T$, we also used the widely accepted values, i.e., $[96,192,336,720]$. For our model, in all configurations, we adopt the GeGLU activation function [16] between the two linear layers in the feed-forward network for our model. Additionally, we use learnable positional embedding parameters for the input data and omit positional embeddings for learnable queries to avoid redundant parameter learning.
For the experiments summarized in Table 2 and Table 9, our model uses three cross-attention layers with embedding size $D=256$, number of attention heads $H=32$. Specifically, to avoid overfitting on small datasets [14], we use patch length 48 on the ETTh1 and ETTh2 datasets. Further details on the hyperparameter settings for these experiments are provided in Table 7.
Table 7: Experimental settings used in Table 2 and Table 9.
@@
Additionally, for the short-term forecasting experiments on the M4 dataset, we employed a slightly different configuration to better suit the nature of short-term time series. The hyperparameter settings for these experiments are detailed in Table 8. The complete results for these short-term experiments are presented in Table 10, while the full results for the long-term forecasting experiments with a fixed 96 input sequence length are provided in the Appendix in Table 9.
Table 8: Experimental settings of M4 Dataset.
Dataset Layers Embedding Size Input Sequence Length Batch Size Epoch Patience Learning Rate
M4-Quartely 3 64 16 32 30 10 $10^{-3}$
M4-Monthly 3 64 36 32 30 10 $10^{-3}$
M4-Yearly 3 64 12 32 30 10 $10^{-3}$
M4-Weekly 3 64 26 32 30 10 $10^{-3}$
M4-Daily 3 64 28 32 30 10 $10^{-3}$
M4-Hourly 3 128 96 32 30 10 $10^{-3}$
Weather 96 0.161 0.207 0.163 0.209 0.186 0.227 0.172 0.220 0.195 0.271 0.198 0.261 0.195 0.236 0.195 0.252 0.266 0.336 0.300 0.384
Electricity 96 0.149 0.237 0.153 0.247 0.190 0.296 0.168 0.272 0.219 0.314 0.180 0.293 0.198 0.274 0.210 0.302 0.201 0.317 0.274 0.368
Traffic 96 0.421 0.270 0.462 0.285 0.526 0.347 0.593 0.321 0.644 0.429 0.577 0.350 0.647 0.384 0.650 0.396 0.613 0.388 0.719 0.391
ETTm1 96 0.318 0.357 0.320 0.357 0.352 0.374 0.338 0.375 0.404 0.426 0.365 0.387 0.353 0.370 0.346 0.374 0.505 0.475 0.672 0.571
ETTm2 96 0.178 0.261 0.175 0.258 0.183 0.270 0.187 0.267 0.287 0.366 0.197 0.296 0.183 0.266 0.193 0.293 0.255 0.339 0.365 0.453
ETTh1 96 0.371 0.395 0.375 0.400 0.460 0.447 0.384 0.402 0.423 0.448 0.426 0.446 0.438 0.433 0.397 0.412 0.449 0.459 0.865 0.713
ETTh2 96 0.287 0.341 0.289 0.341 0.308 0.355 0.340 0.374 0.745 0.584 0.372 0.424 0.322 0.364 0.340 0.394 0.346 0.388 3.755 1.525
Table 10: Full results of univariate short-term forecasting in the M4 dataset. All forecasting horizons are in [6, 48]. The best results are in bold and the second best are underlined.
Models CATS TimeMixer Timesnet PatchTST MICN FiLM DLinear Autoformer Informer
Quarterly SMAPE 9.979 9.996 10.100 10.644 15.214 12.925 12.145 11.338 11.360
MASE 1.164 1.166 1.182 1.278 1.963 1.664 1.520 1.365 1.401
OWA 0.878 0.825 0.890 0.949 1.407 1.193 1.106 1.012 1.027
Monthly SMAPE 12.557 12.605 12.670 13.399 16.943 15.407 13.514 13.958 14.062
MASE 0.916 0.919 0.933 1.031 1.442 1.298 1.037 1.103 1.141
OWA 0.866 0.869 0.878 0.949 1.265 1.144 0.956 1.002 1.024
Yearly SMAPE 13.263 13.206 13.387 16.463 25.022 17.431 16.965 13.974 14.727
MASE 2.967 2.916 2.996 3.967 7.162 4.043 4.283 3.134 3.418
OWA 0.779 0.776 0.786 1.003 1.667 1.042 1.058 0.822 0.881
Others SMAPE 4.560 4.564 4.891 6.558 41.985 7.134 6.709 5.485 24.460
MASE 3.107 3.115 3.302 4.511 62.734 5.09 4.953 3.865 20.960
OWA 0.970 0.982 1.035 1.401 14.313 1.553 1.487 1.187 5.879
Average SMAPE 11.701 11.723 11.829 13.152 19.638 14.863 13.639 12.909 14.086
MASE 1.557 1.559 1.585 1.945 5.947 2.207 2.095 1.771 2.718
OWA 0.838 0.840 0.851 0.998 2.279 1.125 1.051 0.939 1.230
@@ B.1 Performance with Longer Input Sequences
In Table 3, we compared models with the number of parameters, GPU memory consumption, and MSE across different input lengths on ETTm1 with varying input sequence lengths. In this section, we provide comprehensive results on longer input sequence lengths $L=512$. The detailed parameters can be found in Table 11 and the corresponding experimental results are summarized in Table 12. As with unified hyperparameter settings, we follow the settings of the most recent work [21] to ease comparison. Overall, the experimental results clearly illustrate the superiority of CATS over recent forecasting models across multiple datasets and prediction horizons. CATS consistently shows the lowest Mean Squared Error (MSE) and Mean Absolute Error (MAE) across a variety of datasets and forecast horizons. For instance, on Electricity, at the 96 forecast horizon, CATS achieves the best MSE score of 0.144 and the best MAE score of 0.189, underscoring its accuracy in predicting electrical demand.
@@
We provide additional experimental results to support the findings discussed in Section 5.2. Tables 13 and 14 summarize detailed comparisons of the number of parameters, GPU memory consumption, and MSE across different input lengths for the ETTm1 and Weather datasets, respectively. The linear models, TimeMixer and DLinear, exhibit smaller parameters for shorter input lengths. Despite CATS having slightly more parameters than TimeMixer for smaller inputs, it outperforms in terms of memory usage and MSE. This suggests that CATS is more efficient and effective in handling shorter inputs. For PatchTST, the number of parameters does not increase within the actual Transformer backbone as the input length increases. However, due to the need to flatten and project all inputs at the end, the parameters scale linearly with the input length. This highlights a limitation of the Encoder’s architecture. On the other hand, TimeMixer’s parameters grow almost quadratically as the input length doubles. Similarly, DLinear’s parameters increase linearly with the input length. Our proposed model, CATS demonstrates significant efficiency through parameter sharing, where the parameters hardly increase with longer inputs. Notably, from an input length of 336, CATS has fewer parameters than DLinear, showcasing the deep learning model’s advantage in detecting inherent patterns in the data.
@@ B.3 Ablation Study on Query-adaptive Masking
Figure 8: Comparison of performance with query-adaptive masking with two different probabilities, dropout, and using both query-adaptive masking and dropout. The results of $p=0.1$ to $0.7$ indicate a probability masking that is linearly increased proportionally to the horizon predicted by the query.
@@
To push further, we reproduce the experiment of Fig. 7 with other datasets used in forecasting tasks. We illustrate the results of the Weather, Traffic, Electricity, ETTm2, ETTh1, and ETTh2 in Figures 10, 11, 12, 13, 14, and 15, respectively. For each figure, (a) represents the forecasting results, (b) shows the cross-attention score map, and (c) and (d) illustrate the two pairs with the highest attention scores. For all figures, our attention-based explanation successfully discovers similar periodic patterns. Therefore, we believe that our model has the potential to provide a clearer understanding of the mechanisms underlying forecasting predictions. We hope that future research will continue to explore and expand upon this foundation.